# Unsupervised motion segmentation in one go: Smooth long-term model over a video

### v2

@@ Abstract
Human beings have the ability to continuously analyze a video and immediately extract the main motion components. Motion segmentation methods based on deep learning often proceed frame by frame. We want to go beyond this paradigm, and perform the motion segmentation in series of flow fields of any length, up to the complete video sequence. It will be a prominent added value for downstream computer vision tasks, and could provide a pretext criterion for unsupervised video representation learning. In this perspective, we propose a novel long-term spatio-temporal model operating in a totally unsupervised way. It takes as input the volume of consecutive optical flow (OF) fields, and delivers a volume of segments of coherent motion over the video. More specifically, we have designed a transformer-based network, where we leverage a mathematically well-founded framework, the Evidence Lower Bound (ELBO), to infer the loss function. The loss function combines a flow reconstruction term involving spatio-temporal parametric motion models combining, in a novel way, polynomial (quadratic) motion models for the $(x,y)$-spatial dimensions and B-splines for the time dimension of the video sequence, and a regularization term enforcing temporal consistency on the masks. We report experiments on four VOS benchmarks with convincing quantitative results. We also highlight through visual results the key contributions on temporal consistency brought by our method.
@@ 1 Introduction
When dealing with videos, motion segmentation is one of the key issues. Human beings have the ability to continuously analyze a video and immediately extract the main motion components. Computer vision methods in the deep learning era usually proceed frame by frame. We want to go beyond this classical paradigm and segment the different motion entities in series of flow fields of any length, up to the complete video sequence. We thus thoroughly investigate the temporal dimension of the motion segmentation problem. Since the optical flow (OF) yields all the information on the movement between two images of the video sequence, it is natural to base motion segmentation on optical flow. Motion segmentation is a computer vision task in its own, but it is also useful for diverse downstream tasks as independent moving object detection, object tracking, or action recognition, to name a few. It is also leveraged in video object segmentation (VOS), while most often coupled with appearance.
@@
The optical flow field at time $t$ enables to get the segmentation at frame $t$ of the video. However, taking a large time window, and even the whole video sequence, is beneficial, since motion entities are generally consistent throughout a video sequence (or at least within every video shot for long videos). Indeed, temporal coherence is inherent to motion. Therefore, it is essential that motion segmentation takes advantage of it, especially in a long-term perspective.
In this paper, we propose an original holistic method for multiple motion segmentation from optical flow over a video sequence. The main contributions of our work are as follows. Our network takes as input a volume of consecutive optical flows, and delivers consistent motion segmentation maps throughout the video sequence. It involves a transformer module allowing for long-term interactons. It is trained in a completely unsupervised manner, without any manual annotation or ground truth data of any kind. The loss function is inferred by leveraging the Evidence Lower Bound (ELBO) framework, and comprises a flow reconstruction term with original spatio-temporal parametric motion models and an additional term enforcing temporal consistency on the segmentation masks. We model with B-splines the long-term temporal evolution of the motion model parameters. Our method also involves a latent represention of the segment motion augmented with positional embedding.
@@
The rest of the paper is organized as follows. Section 2 describes related work regarding motion segmentation. Section 3 presents our unsupervised network for multiple motion segmentation in one go, embedding long-term temporal consistency. In Section 4, we provide implementation details. Section 5 reports results on four VOS benchmarks with a comparison to state-of-the-art unsupervised motion segmentation methods. Finally, Section 6 contains concluding remarks.
@@ 2 Related work
Motion segmentation aims to break down each frame of a video sequence into components (or segments) of coherent motion. Usually, each motion segment is identified by a motion model, which can be hand-crafted such as affine or quadratic polynomial models or sometimes learned. Motion segmentation has been investigated for decades [19, 23, 31, 40]. However, we will focus on recent methods in this section. Since accurate and efficient methods for estimating optical flow are now available, motion segmentation methods can leverage optical flow as a reliable input. The advent of deep learning methods in computer vision, and more recently the use of transformer-based networks with attention mechanisms [11], has also encompassed motion segmentation.
@@
In [36], a transformer module, more specifically, the slot attention mechanism introduced in [18], is leveraged to perform motion segmentation from the optical flow. As a matter of fact, it addresses a binary segmentation problem, foreground moving object vs background. The loss function is composed of two terms, a flow reconstruction term and an entropy term to make masks as binary as possible. Another approach is adopted in [35] that is able to cope with multiple motion segmentation. Nonlinear subspace filters are learned from stacked deep multi-layer perceptrons. Then, motion segmentation is inferred at inference by applying $K$-means to the output embeddings. In [21], the Expectation-Maximization (EM) framework is leveraged to define the loss function and the training procedure for an unsupervised frame-by-frame motion segmentation, where 12-parameter quadratic motion models are involved. The same authors also proposed an extension of that method in [4] to better handle the temporal dimension of the problem, by considering triplets of flows as input. In [37], the authors developed an adversarial method whose aim is to generate a mask hiding the input optical flow, where an inpainter network attempts to recover the flow within the mask. The rationale is that no correct reconstruction of the inside flow can be achieved from the outside flow, if the hidden area exhibits an independent motion and then constitutes a motion segment.
@@
The temporal dimension of motion segmentation has been considered in various ways [31]. Regarding deep learning approaches, motion segmentation at time $t$ is improved at training time in [36], by taking, as input, flows between $t$ and several time instants before and after $t$. The authors of [8] consider several consecutive RGB frames as input of their self-supervised method. Optical flow is only computed at training time, and the loss function also comprises a temporal consistency term. However, the latter is not applied to two consecutive segmentation masks, but for pairings between the same frame $t$ and another (more or less distant) one. In [9], spatio-temporal transformers were designed for video object segmentation involving temporal feature propagation.
@@
VOS usually requires motion segmentation, often coupled with appearance [5, 10, 38]. It was addressed with supervised or semi-supervised methods as in [7, 9, 12], but also with unsupervised methods [37, 36, 21]. VOS is concerned with the segmentation of primary objects, i.e., object moving in the foreground of a scene and possibly tracked by the camera [41]. Thus, VOS generally involves binary ground truth segregating primary moving object against background. This is for instance the case for the DAVIS2016 benchmark [26]. Recent works have revisited the way of coupling appearance and motion for VOS. The AMD method [17] includes two pathways, the appearance one and the motion one. If it does not use optical flow as input and brings out the objectness concept, it nevertheless relies on a coarse motion representation. In [39], the authors introduce deformable sprites with a video auto-encoder model (optimized at test time), to segment scene elements through the video, by leveraging texture and non-rigid deformations. The work [27] leverages pretrained appearance features (using DINO) and optical flow. The objective function is derived from a form of spectral clustering and applied to the entire video. However, optimization is required at test time. The RCF method [16] involves learnable motion models, and is structured in two stages, a motion-supervised object discovery stage, and a refinement stage with residual motion prediction and high-level appearance supervision. However, the method cannot distinguish objects undergoing different motions. In [6], the prediction of motion patterns is used at training time to learn objectness from videos. The DivA method [2] promotes divided attention. It is based on the same principle as in [37] that motion segments are mutually uninformative, but it is not limited to binary segmentation. It can segment a variable number of moving objects from optical flow, by leveraging a slot attention mechanism guided by the image content through a cross-modal conditional slot decoder.
Our fully unsupervised approach differs from these previous works in several respects. We rely only on optical flow and take a volume of OF fields as input, providing a volume of consistent segmentation maps. We introduce B-splines to define parametric motion models able to correctly handle motion evolution over time, and we express long-term temporal consistency. We infer the loss function from the Evidence Lower Bound (ELBO) framework. In [4], the authors also deal with temporal consistency. However, they only took a triplet of optical flow fields as input, they did not resort to transformers, and the temporal part of the motion model was just a linear (first-order polynomial) model in time. In addition, their temporal linkage of the consecutive motion segmentations over a video sequence was an ad hoc post-processing. In contrast, we define a fully integrated long-term model that is end-to-end trained in a unsupervisedly way. We will call our method LT-MS method, for long-term motion segmentation, and in contrast, we will refer to the method [4] as the ST-MS method, for short-term motion segmentation.
@@ 3 Long-term Motion Segmentation Method
We have designed a transformer-based network for multiple motion segmentation from optical flow. It is inspired from the MaskFormer architecture [4], but it only comprises one head corresponding to the mask prediction, as described in Fig.1. The network takes as input a volume, of flexible temporal length, comprising several consecutive optical flow fields. Temporal consistency is expressed in two main ways at the training stage. Firstly, we associate a space-time motion model with each segment to characterize its motion along with the evolution of the latter over time. Secondly, the loss function comprises an additional term enforcing consistent labeling of the motion segments over the volume. Adding the transformer to the latent space of the Unet allows for interaction between distant features in the input, and thus for long-term segmentation. We only incorporate a transformer decoder on the downsampled latent space of the 3D U-net, which induces a limited additional computational load.
@@ 3.1 Spatio-temporal parametric motion model
The set of $T$ consecutive flows will be designated as a space-time volume (or volume, to make it short). The volume could even be the whole video sequence. Our space-time motion model is estimated through B-spline functions [33] for its temporal dimension. More precisely, we consider a B-spline function of order $n$ in the variable $t$ over the space-time volume. For the spatial extension, we assign a parametric motion model $\tilde{f}_{\theta_{k}}$ to each motion segment $k,k=1,..,K$, where $\theta_{k}$ specifies the spatial motion model $\tilde{f}$ for segment $k$. We take $n=3$. The number $L$ of control points is given by $L=2+\lfloor\frac{T-2}{\nu}\rfloor$, where $\nu$ manages the temporal frequency of control points. We put a control point at both ends of the volume ($t=1$, $t=T$), and the other control points are equidistantly placed between the two. The control points are often not located at time points of the video.
@@
The space-time spline-based motion model is illustrated in Fig.2. Just for this illustration, the motion models are computed within the two segments, foreground moving object and background, provided by the ground truth. The estimated motion model for the foreground moving object is able to capture the periodic nature of the swing motion as demonstrated by the plots of the computed motion model parameters. Also, the motion model computed in the background segment perfectly fits the camera motion. Articulated motion (woman’s legs) would require multiple-motion segmentation.
@@
Figure 2: Illustration of the spatiotemporal spline-based motion model. Top row: input flows displayed with the HSV code for the swing video of DAVIS2016, binary segmentation ground truth, flows generated by the estimated spline-based motion models for the two segments. Bottom row: plot of the temporal evolution of the six estimated motion model parameters corresponding to the horizontal flow component for the foreground moving object. The $x$-axis is the video frame index.
@@
Any parametric motion model could be considered. We take a polynomial expansion in the coordinates $x$ and $y$ of any point in the image, with one polynomial for each component of the 2D flow. The use of polynomial is a usual choice for 2D motion modelling, since it is easy to estimate and enables a straightforward physical interpretation. More specifically, we select the 12-parameter quadratic motion model to be able to account for continuously varying depth surface of the objects in the scene, especially for the whole background, and for complex object or camera motions. In contrast, the affine and the 8-parameter quadratic motion models assume a planar object surface. Indeed, the latter exactly corresponds to the projection in the image of the 3D rigid motion of a planar surface. It is equivalent for velocity fields to the homography transform for points. However, in presence of severe depth effects (strong depth discontinuities) and camera motion, the static scene cannot be represented by a single motion model due to motion parallax produced by static objects located in the foreground. Regarding first the spatial dimension of the motion model, the 2D flow vector yielded by the full quadratic motion model at point $(x,y)$ writes:
$\tilde{f}_{\theta}(x,y)=(\theta_{1}+\theta_{2}x+\theta_{3}y+\theta_{7}x^{2}+\theta_{8}xy+\theta_{9}y^{2},\theta_{4}+\theta_{5}x+\theta_{6}y+\theta_{10}x^{2}+\theta_{11}xy+\theta_{12}y^{2})^{T}.$ (1)
@@
B-Splines are modelling the evolution of the parameters of the motion model over time. The resulting motion model at a given time is then a weighted average of the close control ones (each control one being a quadratic model as in eq.(1). By denoting $S_{n}(\theta_{k})$ the subscript of $\tilde{f}$, we emphasize that the motion model parameters for each segment $k$ are estimated through the B-spline functions:
@@
where $B_{n,l}$ is the $l^{th}$ B-spline, $\theta_{k,l}$ corresponds to the $l^{th}$ control point of the spline, and $\tilde{f}_{\theta_{k,l}}$ is defined by eq.(1).

### v3

# Segmenting the motion components of a video: A long-term unsupervised model

@@ Abstract
Human beings have the ability to continuously analyze a video and immediately extract the motion components. We want to adopt this paradigm to provide a coherent and stable motion segmentation over the video sequence. In this perspective, we propose a novel long-term spatio-temporal model operating in a totally unsupervised way. It takes as input the volume of consecutive optical flow (OF) fields, and delivers a volume of segments of coherent motion over the video. More specifically, we have designed a transformer-based network, where we leverage a mathematically well-founded framework, the Evidence Lower Bound (ELBO), to derive the loss function. The loss function combines a flow reconstruction term involving spatio-temporal parametric motion models combining, in a novel way, polynomial (quadratic) motion models for the spatial dimensions and B-splines for the time dimension of the video sequence, and a regularization term enforcing temporal consistency on the segments. We report experiments on four VOS benchmarks, demonstrating competitive quantitative results, while performing motion segmentation on a whole sequence in one go. We also highlight through visual results the key contributions on temporal consistency brought by our method.
@@ 1 Introduction
When dealing with videos, motion segmentation is one of the key issues. Human beings have the ability to continuously analyze a video and immediately extract the diverse motion components. We want to design a similar paradigm to provide a coherent and stable motion segmentation over the video sequence. Since optical flow (OF) yields all the information on the movement between two images of the video sequence, it is natural to base motion segmentation on optical flow. Motion segmentation is a computer vision task in its own, but it is also useful for downstream tasks. Indeed, it is essential to have an accurate and reliable motion segmentation for independent moving object detection, multiple object tracking, or video understanding, to name a few. Motion segmentation based on optic flow has recently received renewed attention in the context of deep learning [29, 47, 49, 50], but the temporal extension remains to be further investigated.
@@
In this paper, we propose an original long-term unsupervised model for segmenting all the motion components of a video, while taking optical flow as input. It must therefore be considered first and foremost as an unsupervised method performing optical flow segmentation (OFS) over an entire sequence. We assume that optical flow is readily available, computed by one of the recent efficient optical flow methods [43]. The optical flow field at time $t$ enables to get the motion segmentation map at frame $t$ of the video. However, taking a large time window, and even the whole video sequence, is beneficial, since motion entities are generally consistent throughout a video sequence (or at least within every video shot for long videos). Indeed, temporal coherence is inherent to motion.
Motion components (or segments) observed in an image sequence may come from a variety of sources. The first that springs to mind is the case of one or several independent moving objects in the viewed scene. In case of articulated motion, a motion segment will ensue from each part of the articulated body. If the camera is itself moving, the static scene exhibits an apparent motion in the video, which may induce several motion segments according to the depth and/or orientation of the static objects in the scene. In particular, a static object in the foreground of the scene will have an apparent motion different than the scene background one, which is usually referred to as motion parallax [15].
Even if we take an object-oriented view of video content, motion appears as an intrinsic, globalizing property of an object. It is more immediate to segment an object based on its motion than on its appearance. Indeed, motion has been used in the training stage to facilitate image segmentation at inference time [7], or as a criterion to validate image segments [25]. Motion segmentation is also frequently involved in video object segmentation (VOS).
Admittedly, better performance is obtained in VOS, if appearance and motion are combined, compared to the single use of motion. However, there is a bias in the VOS datasets: the primary object to be detected is not only salient by its motion compared to the background, but also by its appearance and even its location, the object being generally placed in the image center. Yet, combining the two modalities may confine their use to videos where the appearance of the object of interest can be easily represented by deep features, like DINO features [4] that are trained on a huge set of natural images from the Internet. Indeed, such algorithms could not be easily applied to videos depicting other types of objects and contents, such as satellite images, biological images, underwater or industrial videos. Models trained on optical flow have a much better generalization and domain transfer ability. Moreover, relying on motion allows to deal with contents where appearance is not discriminating, for instance for crowd scenes [28], or for moving camouflaged animals in natural scenes [18].
Our long-term unsupervised spatio-temporal model is a significant extension of the model we introduced in [2]. For the sake of clarity, from now on, we will name the method presented in this paper LT-MS (for long-term motion segmentation) and the previous one described in [2] ST-MS (for short-term motion segmentation).
First, we have designed a fully end-to-end method, whereas ST-MS required an ad-hoc post-processing step to achieve the temporal linkage beyond three successive frames. This was made possible by defining a new space-time parametric motion model to handle more elaborated temporal evolutions. The spatial part of the motion model remains a polynomial model (e.g., a quadratic one), while the temporal part results from a B-spline approximation. Besides, we have added a transformer (only the decoder part) on the latent space of the 3D U-net, allowing for long-range interactions. While ST-MS takes only three consecutive flows as input, LT-MS can accommodate at inference time any length of series of flow fields, ten, twenty, one hundred, or even the whole sequence. Second, the LT-MS loss function is similar to the ST-MS one, but it is now inferred in a mathematical well-founded manner, leveraging the Evidence Lower Bound (ELBO) framework. We also propose a mathematical development to motivate the temporal consistency term of the loss function. Third, we have added a fourth dataset in the experimental evaluation, the DAVIS2017-motion dataset introduced in [47].
@@
The rest of the paper is organized as follows. Section 2 describes related work regarding motion segmentation. Section 3 presents our unsupervised network for long-term multiple motion segmentation. In Section 4, we provide implementation details. Section 5 reports results on four VOS benchmarks with a comparison to state-of-the-art unsupervised motion segmentation methods. Finally, Section 6 contains concluding remarks.
@@ 2 Related work
Motion segmentation aims to break down each frame of a video sequence into components (or segments) of coherent motion. Usually, each motion segment is identified by a motion model, which can be hand-crafted such as affine or quadratic polynomial models, or sometimes learned [42]. Motion segmentation has been investigated for decades [8, 26, 31, 40, 45, 46, 53]. However, we will focus on recent methods in this section. We refer the reader to [29, 31, 53] for a review of older works. Since accurate and efficient methods for estimating optical flow are now available, motion segmentation methods can leverage optical flow as a reliable input. The advent of deep learning, and more recently the use of transformers with attention mechanisms [14], has also encompassed motion segmentation.
@@
In [49], a transformer module, more specifically, the slot attention mechanism introduced in [24], is leveraged to perform motion segmentation from optical flow. As a matter of fact, it addresses a binary segmentation problem, foreground moving object vs background. The loss function is composed of two terms, a flow reconstruction term and an entropy term to make masks as binary as possible. In [2], it is assumed that camera rotation can be removed in the optical flow field, which leads to a simplified scheme to detect independently moving objects in a self-supervised way. Another approach is adopted in [48] that is able to cope with multiple motion segmentation. Nonlinear subspace filters are learned from stacked deep multi-layer perceptrons. Then, motion segmentation is inferred at inference by applying $K$-means to the output embeddings. In [29], we leveraged the Expectation-Maximization (EM) framework to define the loss function and the training procedure for an unsupervised frame-by-frame motion segmentation, where 12-parameter quadratic motion models are involved. In [50], the authors developed an adversarial method whose aim is to generate a mask hiding the input optical flow, where an inpainter network attempts to recover the flow within the mask. The rationale is that no correct reconstruction of the inside flow can be achieved from the outside flow, if the hidden area exhibits an independent motion and then constitutes a motion segment.
The temporal dimension of motion segmentation has already been considered in different ways. Regarding classical approaches, in [33] the motion partition at time $t$ was predicted from the one obtained at time $t-1$ using the affine motion models estimated between images $t-1$ and $t$ for each segment, within a robust MRF-based method. The authors in [40] showed that it was beneficial to introduce temporal layer constancy over several frames to perform motion segmentation in a MRF-based and graph-cut optimization framework. In [32], large time windows were taken into account, but point trajectories were used as input within a spectral clustering method.
@@
Regarding deep learning approaches, motion segmentation at time $t$ is improved at training time in [49], by taking, as input, flows between $t$ and several time instants before and after $t$. The authors of [10] consider several consecutive RGB frames as input of their self-supervised method. Optical flow is only computed at training time, and the loss function also comprises a temporal consistency term. However, the latter is not applied to two consecutive segmentation masks, but for pairings between the same frame $t$ and another (more or less distant) one. In [11], spatio-temporal transformers were designed for video object segmentation involving temporal feature propagation. In [52], the authors introduce deformable sprites with a video auto-encoder model, but optimized at test time, to segment scene elements through the video. It leverages texture and non-rigid deformations. The work [36] involves pretrained appearance features (using DINO) and optical flow. The objective function is derived from a form of spectral clustering and applied to the entire video. However, optimization is required at test time. In [2], we explicitly handled the temporal dimension of the problem by considering triplets of consecutive flows as input, a space-time polynomial motion model, and a temporal consistency term in the loss function.
@@
Since motion segmentation can be used in the context of VOS and that we will comparatively evaluate our method on VOS benchmarks, we briefly review VOS methods. The latter often combine appearance and motion features [6, 12]. It was addressed with supervised or semi-supervised methods as in [9, 11, 16, 51], but also with unsupervised methods [29, 49, 50]. VOS is concerned with the segmentation of primary objects, i.e., object moving in the foreground of a scene and possibly tracked by the camera [54]. Thus, VOS generally involves binary ground truth segregating primary moving object against background. This is for instance the case for the DAVIS2016 benchmark [35]. Recent works have revisited the way of coupling appearance and motion for VOS. The AMD method [23] includes two pathways, the appearance one and the motion one. If it does not use optical flow as input and brings out the objectness concept, it nevertheless relies on a coarse motion representation. The RCF method [21] involves learnable motion models, and is structured in two stages, a motion-supervised object discovery stage, and a refinement stage with residual motion prediction and high-level appearance supervision. However, the method cannot distinguish objects undergoing different motions. In [7], the prediction of motion patterns is used at training time to learn objectness from videos. The DivA method [1] promotes divided attention. It is based on the same principle as in [50] that motion segments are mutually uninformative, but it is not limited to binary segmentation. It can segment a variable number of moving objects from optical flow, by leveraging a slot attention mechanism guided by the image content through a cross-modal conditional slot decoder.
Our fully unsupervised long-term approach differs from these previous works in several respects as outlined in the introduction.
@@ 3 Long-term Motion Segmentation Method
We have designed a transformer-based network for multiple motion segmentation from optical flow. It is inspired from the MaskFormer architecture [5], but it only comprises one head corresponding to the mask prediction, as described in Fig.1. The network takes as input a volume, of flexible temporal length, comprising several consecutive optical flow fields. Temporal consistency is expressed in two main ways at the training stage. Firstly, we associate a space-time motion model with each segment to characterize its motion along with the evolution of the latter over time. Secondly, the loss function comprises an additional term enforcing consistent labeling of the motion segments over the volume. Adding the transformer to the latent space of the 3D U-net allows for interaction between distant features in the input, and thus for long-term segmentation. We only incorporate a transformer decoder on the downsampled latent space, which induces a limited additional computational load.
@@ 3.1 Spatio-temporal parametric motion model
The set of $T$ consecutive flows will be designated as a space-time volume (or volume, to make it short). The volume could even be the whole video sequence. Our space-time motion model is estimated through B-spline functions [44] for its temporal dimension. More precisely, we consider a B-spline function of order $n$ in the variable $t$ over the space-time volume. For the spatial extension, we assign a parametric motion model $\tilde{f}_{\theta_{k}}$ to each motion segment $k,k=1,..,K$, where $\theta_{k}$ specifies the spatial motion model $\tilde{f}$ for segment $k$. In practice, we take $n=3$, which proved sufficient to handle motion evolution. The number $L$ of control points is given by $L=2+\lfloor\frac{T-2}{\nu}\rfloor$, where $\nu$ manages the temporal frequency of control points. We put a control point at both ends of the volume ($t=1$, $t=T$), and the other control points are equidistantly placed between the two. The control points are often not located at time points of the video.
@@
The space-time spline-based motion model is illustrated in Fig.2. Just for this illustration and for the sake of clarity, the motion models are computed within the two segments, foreground moving object and background, provided by the ground truth. The estimated motion model for the foreground moving object is able to capture the periodic nature of the swing motion as demonstrated by the plots of the computed motion model parameters. Also, the motion model computed in the background segment perfectly fits the camera motion. Articulated motion (woman’s legs) would require multiple-motion segmentation.
@@
Figure 2: Illustration of the spatiotemporal spline-based motion model. Top group: respectively, input flows displayed with the HSV code for the swing video of DAVIS2016, binary segmentation ground truth, flows generated by the estimated spline-based motion models for the two segments. Bottom row: plot of the temporal evolution of the six estimated motion model parameters corresponding to the horizontal flow component for the foreground moving object. The $x$-axis is the video frame index.
@@
Any parametric motion model could be considered. We take a polynomial expansion in the coordinates $x$ and $y$ of any point in the image, with one polynomial for each component of the 2D flow. The use of polynomial is a usual choice for 2D motion modelling [31], since it is easy to estimate and enables a straightforward physical interpretation. More specifically, we select the 12-parameter quadratic motion model to be able to account for continuously varying depth surface of the objects in the scene, especially for the whole background, and for complex object or camera motions. In contrast, the affine and the 8-parameter quadratic motion models assume a planar object surface. Indeed, the latter exactly corresponds to the projection in the image of the 3D rigid motion of a planar surface. It is equivalent for velocity fields to the homography transform for points. However, in presence of severe depth effects (strong depth discontinuities) and camera motion, the static scene cannot be represented by a single motion model due to motion parallax produced by static objects located in the foreground.
Regarding first the spatial dimension of the motion model, the 2D flow vector yielded by the full quadratic motion model at point $(x,y)$ writes:
$\tilde{f}_{\theta}(x,y)=$ $(\theta_{1}+\theta_{2}x+\theta_{3}y+\theta_{7}x^{2}+\theta_{8}xy+\theta_{9}y^{2},$
$\theta_{4}+\theta_{5}x+\theta_{6}y+\theta_{10}x^{2}+\theta_{11}xy+\theta_{12}y^{2})^{T}.$ (1)
@@
B-Splines are modelling the evolution of the parameters of the motion model over time. The resulting motion model at a given time is then a weighted average of the close control ones, each control one being a quadratic model as in eq.(3.1). By denoting now $S_{n}(\theta_{k})$ the subscript of $\tilde{f}$, we emphasize that the motion model parameters for each segment $k$ are estimated through the B-spline functions:
@@
where $B_{n,l}$ is the $l^{th}$ B-spline, $\theta_{k,l}$ corresponds to the $l^{th}$ control point of the spline, and $\tilde{f}_{\theta_{k,l}}$ is defined by eq.(3.1).